Station BORG, Borgarfjordur, Asbjarnarstadir, Iceland

BORG operational status:


IDA engineers visited the station and installed a new Trillium 360 seismometer with modifications to improve performance , upgraded power to the equipment rack, and installed an EdgeRouter to improve network communications.


IDA engineers are onsite to replace the KS54000 seismometer with a Trillium360 seismometer, and install a new Episensor strong motion sensor, Hyperion infrasound sensor, weather station, and new station computers.


The station operator replaced the FBA-23 with an episensor strong motion sensor on day 2016,256-14:36. However, the Episensor is defective and will also need to be replaced.


The station operator removed the steel plate beneath the STS-2.5 seismometer and re-installed the seismometer on day 2012,172-15:48. Removing the plate improved data quality for the horizontal channels from the STS-2.5 seismometer.


The station operator replaced the STS-2 with a new STS-2.5 seismometer on day 2012,108. The data for the new sensor was non-seismic until the masses were re-centered on day 2012,109.


IDA station BORG has been upgraded to a Quanterra Q330HR system. Sensors include KS54000 and STS-2 seismometers, and an FBA-23 accelerometer.


The station was down from day 2010,154 to 2010,160 due to a problem with the DAS. Data acquistion was restored after the station operator cycled power to the DAS.


A new GPS cable was installed on day 2008,067 and resolved the GPS clock problem.


The timing for data from IDA station BORG was unreliable from about day 2008,055 to 2008,060 due to problems with the GPS clock. A new cable for the GPS clock is being shipped to the station.


The timing for data from IDA station BORG was unreliable from about day 2008,042 to 2008,048 due to problems with the GPS clock.


The station was down as of about day 2008,032-10:10 due to problems with the GPS clock. Some wiring to the GPS clock antenna was damaged by horses.


An IDA engineer visited the station and replaced the GPS clock since the GPS clock sent to the station failed.


The station operator replaced a faulty GPS clock on day 304. The GPS clock had been in an unlocked state since day 228. Timing of data during this period is uncertain. Data are continuous and data waveform quality remained unaffected.


Station operator reports replacing the ISP and UPS for the VSAT (05/11). Telemetry to the station is down since 05/11.


Telemetry restored on 8/10/2004.


Telemetry down as of 7/28/2004. Station remains operational.


The data acquisition system has been upgraded to a MK8 system. Additional sensors have been added including an FBA-23 accelerometer and a Paroscientific microbarograph.


DAT drive and on-site spares have failed. Replacements shipped 2003.03.17.


The operator installed the DAT drive shipped from IDA and restored data acquisition.


The operator notified IDA staff that the DAT drive has failed and that there are no spares on site.


New EPROMs installed to correct DSP problem.


Station is operating normally.

Telemetry to the station is not available due to difficulties communicating through the firewall at IMO. Alternate communication routes are being pursued.